# Efficient sequential Bayesian inference for state-space epidemic models using ensemble data assimilation

### v1

@@ 2.2 Ensemble Kalman Filter
$\widehat{p}_{\text{enkf}}^{N_{x}}(y_{t}|y_{1:t-1},\theta)=\widehat{\mathcal{N}}_{\text{unb}}\left(y_{t};\penalty 10000\ H\widehat{\mu}_{t|t-1},\penalty 10000\ H\widehat{\Sigma}_{t|t-1}H^{\top}+V_{t}\right).$ (17)
@@
Unlike in the case of the particle filter (Andrieu et al. 2010), the resulting posterior approximation $\widehat{p}_{\text{enkf}}^{N_{x}}(\theta|y_{1:t})\propto\widehat{p}_{\text{enkf}}^{N_{x}}(y_{1:t}|\theta)\,p(\theta)$, still does not target the true posterior (due to the Gaussian assumption). However, when the Gaussian assumption holds, the resulting posterior no longer depends on $N_{x}$, eliminating the finite-ensemble bias and yielding a stable target for Bayesian inference (Ghurye & Olkin 1969, Drovandi et al. 2022). The EnKF procedure, including unbiased likelihood estimation, is outlined in Algorithm 1. This estimator can then be incorporated into a Metropolis–Hastings algorithm to target the sequence of posterior distributions $\{p(\theta),\penalty 10000\ p(\theta|y_{1}),\penalty 10000\ p(\theta|y_{1:2}),\ldots,\penalty 10000\ p(\theta|y_{1:T})\}$.
@@
Inputs: Observation $y_{1:T}$, Number of ensemble members $N_{x}$, Inital state distribution $p(x_{0})$, Parameter vector $\theta$.
@@
4:  Compute the observation variance, ensemble forecast mean, and variance: $V_{t}$, $\widehat{\mu}_{t|t-1}$, and $\widehat{\Sigma}_{t|t-1}$.
@@
$\widehat{p}_{\text{enkf}}^{N_{x}}(y_{t}|y_{1:t-1},\theta)=\widehat{\mathcal{N}}_{\text{unb}}\left(y_{t};\penalty 10000\ H\widehat{\mu}_{t|t-1},\penalty 10000\ H\widehat{\Sigma}_{t|t-1}H^{\top}+V_{t}\right).$
@@ 2.3 Ensemble SMC2
The EnKF algorithm described previously assumes that the model parameters $\theta$ are known. In practice, these parameters are rarely known, and it is preferable to assign a prior distribution $p(\theta)$ that reflects prior knowledge. To account for this uncertainty, inference on $\theta$ can be performed by embedding the EnKF within an SMC sampler over the parameter space (Del Moral et al. 2006). This approach follows the general principle of SMC2 (Chopin et al. 2013), which traditionally employs a particle filter to estimate the likelihood. At each observation time $t$, the posterior distribution of the parameters given data $y_{1:t}$ satisfies the recursive relation
@@
As the incremental likelihood $p(y_{t}|y_{1:t-1},\theta)$ is intractable in a nonlinear system, at time $t$, the weight of the $\theta$-particle $\theta^{m}$ is updated using an estimator $\widehat{p}_{\text{enkf}}^{N_{x}}(y_{t}|y_{1:t-1},\theta^{m})$ obtained from the associated EnKF. This differs from related work of Wu et al. (2022), where the EnKF is used to construct proposal kernels in the SMC sampler but not as a likelihood estimator. As in standard SMC2, particle degeneracy may occur over time. To mitigate this, resampling (described in Algorithm B5) is triggered whenever the effective sample size (ESS) drops below $N_{\theta}/2$ (Chopin et al. 2013, 2020). This is followed by a jittering mechanism that rejuvenates the particles by proposing new candidates, which are then accepted or rejected through a few iterations of Particle Marginal Metropolis–Hastings (PMMH). Since the weighted particles already approximate $p(\theta|y_{1:t})$ (under the Gaussian assumption), applying a Metropolis-Hastings update to each one can only enhance this approximation (see Algorithm 2).
@@
$\vartheta(\theta^{*},\theta)=\min\left\{1,\penalty 10000\ \dfrac{\widehat{p}_{\text{enkf}}^{N_{x}}(y_{1:t}|\theta^{*})p(\theta^{*})}{\widehat{p}_{\text{enkf}}^{N_{x}}(y_{1:t}|\theta)p(\theta)}\times\dfrac{q(\theta|\theta^{*})}{q(\theta^{*}|\theta)}\right\}$
@@
The eSMC2 algorithm can be viewed as a variant of the standard SMC2, where the particle filter–based likelihood is replaced by an EnKF approximation in (18). From a practical perspective, this modification is expected to have only a slight effect on the posterior estimates of $\theta$ and $x_{1:t}$. The complete procedure is detailed in Algorithm 3, with schematic overview in Figure 2.
To propagate parameter uncertainty in the hidden state estimation, we approximate the marginal posterior distribution of the latent states at time $t$, $p(x_{t}|y_{1:t})$, by integrating over a subset of parameter particles, $\{\theta^{j}\}_{j=1}^{N_{c}}\sim p(\theta|y_{1:T})$, drawn from Algorithm 3 output (Steyn et al. 2025). In this work, we set $N_{c}=100$, which in our numerical examples (Section 3) provides satisfactory coverage of the posterior. For each sampled $\theta^{j}$, an EnKF with $N_{x}$ ensemble members is run to approximate the filtering distribution $p(x_{t}|y_{1:t},\theta^{j})$, and the resulting trajectories are combined across all parameter samples to form an empirical approximation of the state posterior:
@@
It is important to note a key theoretical distinction between SMC2 and eSMC2. Standard SMC2 is “exact”: as the number of state and parameter particles $N_{x},N_{\theta}\to\infty$, it converges to the true posterior $p(\theta|y_{1:t})$. In contrast, eSMC2 is generally biased because the EnKF approximates the incremental likelihood $p(y_{t}|y_{1:t-1},\theta)$ via a Gaussian assumption. Nevertheless, the EnKF likelihood typically has low variance (Katzfuss et al. 2020), and unbiased corrections are possible under mild conditions (Drovandi et al. 2022). Empirical studies show that EnKF-based likelihoods yield accurate and stable inference for both linear and moderately nonlinear SSMs (Khalil et al. 2015, Drovandi et al. 2022). Thus, while not formally exact, eSMC2 provides practically accurate posterior estimates at much lower computational cost.
@@ 3.1 Inference on a simulated epidemic
To assess the performance of the proposed eSMC2 algorithm, we first conducted two simulation studies using the SEIR model from Section 2. All algorithms were implemented in Python and executed on a desktop computer equipped with an Intel Core i7-1300H processor (3.40 GHz). Unless otherwise stated, simulations assumed a population size of N=500, [^0] N=500{,}000 with initial conditions $I_{0}=10$, $S_{0}=N-I_{0}$, and $E_{0}=R_{0}=0$. Daily incidence $y_{t}$ was generated according to the Poisson observation model in Eq. (3), assuming perfect case reporting ($\rho=1$).
The two examples differ primarily in the temporal structure of the transmission rate $\beta_{t}$ and the number of observation points. Example 1 depicts a moderately time-varying epidemic with an initial surge followed by a mild resurgence, capturing short-term fluctuations in transmission. Example 2, by contrast, exhibits a smoother and more prolonged pattern that sustains infections for longer and decays more gradually. The latter setting was designed to emulate the dynamic observed in the 2022 U.S. mpox epidemic, where behavioral changes and intermittent control measures led to prolonged and uneven declines in case counts.
@@
Synthetic data were generated over $T=60$ time points with time-varying transmission rate $\beta_{t}=0.3\cdot\exp\big(\sin(2\pi t/55)-t/80\big)$. The incubation and infectious periods were set to $\alpha=1/2$ and $\gamma=1/7$, respectively. Each state ensemble was initialized using truncated normal distributions: $S_{0}\sim\mathcal{TN}_{[0,\,\infty]}(5\times 10^{5},0.2^{2})$, $I_{0}\sim\mathcal{TN}_{[0,\,\infty]}(10,0.2^{2})$ and $E_{0}=R_{0}=0$ .The priors were specified as $\beta_{0}\sim\mathcal{N}(0.3,0.01^{2}),\penalty 10000\ \alpha\sim\mathcal{TN}_{[0,\,\infty]}(0.6,0.3^{2}),\penalty 10000\ \gamma\sim\mathcal{TN}_{[0,\,\infty]}(0.2,0.1^{2}),\penalty 10000\ \nu_{\beta}\sim\mathcal{U}(0,0.5).$
@@
Synthetic data were generated over $T=100$ time points with a time-varying transmission rate $\beta_{t}=0.5\cdot\exp\Big(-(t-15)^{2}/20^{2}\Big)+0.065$, and fixed parameters $\alpha=1/3$ and $\gamma=1/8$. The priors were specified as $\beta_{0}\sim\mathcal{N}(0.35,0.01^{2}),\penalty 10000\ \alpha\sim\mathcal{TN}_{[0,\,\infty]}(0.4,0.2^{2}),\penalty 10000\ \gamma\sim\mathcal{TN}_{[0,\,\infty]}(0.12,0.2^{2}),\penalty 10000\ \nu_{\beta}\sim\mathcal{U}(0,0.3).$ Initial state ensemble members were identical to Example 1.
For both simulation studies, we compared the proposed eSMC2 algorithm with the standard SMC2 implementation that employs a bootstrap particle filter (BPF) in its inner layer (see Appendix B). In both cases, we used $N_{\theta}=1000$ parameter particles and $N_{x}=200$ ensemble members (or particles). For simplicity and comparability, the value of $N_{x}$ was kept fixed throughout the timeline. Parameter rejuvenation was triggered when the effective sample size of the parameter particles dropped below 50% of $N_{\theta}$, followed by five iterations of the PMMH kernel.
Filtered estimates in Figures 3 and 4 correspond to a single representative run of each method. The results illustrate that both eSMC2 (blue) and SMC2 (orange) accurately capture the epidemic dynamics in Examples 1 and 2. For both methods, the predicted incidence closely follows the observed data, with 95% credible intervals consistently encompassing the actual case counts, indicating reliable uncertainty quantification. Estimates of the time-varying transmission rate $\beta_{t}$ and the effective reproduction number $R_{\text{eff}}(t)$ are broadly consistent across methods and closely align with the ground truth, demonstrating that both approaches yield comparable assessments of the evolving transmission dynamics. Figures C.1-C.2 in Appendix C illustrate that the filtering estimates closely track the true latent trajectories, with credible intervals providing appropriate coverage.
@@
Figures 5 and 6 illustrate the temporal evolution of the posterior means of the inferred parameters, along with their marginal posterior densities at the final time step, each summarizing five independent runs. Across runs, both SMC2 and eSMC2 accurately recover the ground truth. For the latency and recovery rates ($\alpha$ and $\gamma$), the posterior means obtained from both algorithms are closely aligned, with credible intervals narrowing steadily as more data are assimilated, indicating progressive information gain. The only notable exception is the diffusion parameter $\nu_{\beta}$ in Example 2, where a slightly larger discrepancy is observed. This difference likely arises from the stochastic nature of the log-transmission process combined with the approximation introduced by the EnKF-based likelihood in eSMC2. Importantly, these differences do not introduce systematic bias or distort the inferred $\beta_{t}$ trajectory. The joint posterior contour plots (Figures C.3-C.4, Appendix C) further reveal strong agreement in posterior geometry and dependence structure: the contours for SMC2 and eSMC2 largely overlap, are centered near the true values, and display similar correlation patterns among parameters. The only visible difference appears along the $\nu_{\beta}$ dimension. Additional sensitivity analyses, reported in Appendix D examines the effect of the number of state particles ($N_{x}$) and variations in the reporting fraction ($\rho$), confirming that the posterior inferences are robust to ensemble size and show only moderate changes when $\rho$ is correctly specified. The analyses also investigate the impact of non-informative (flat) priors, demonstrating that posterior estimates remain stable under weakly informative prior specifications.
@@
Results in Table 1 summarize the accuracy and computational cost of each method, averaged over five independent runs. The mean absolute error (MAE) and root mean squared error (RMSE) quantify discrepancies between the filtering means and the ground truth for both selected latent state ($Z_{t},\penalty 10000\ \beta_{t}$) and static parameters ($\alpha,\penalty 10000\ \gamma$). Across both examples, eSMC2 achieves error magnitudes closely matching those of SMC2, indicating that the ensemble approximation preserves inferential accuracy. In terms of computational effort, eSMC2 offers a substantial advantage. In Example 1, the average CPU time is approximately 467 seconds for eSMC2, compared with 2760 seconds for SMC2, corresponding to a reduction by a factor of about 4.87. Similarly, in Example 2, eSMC2 reduces average runtime to 639 seconds versus 3929 seconds for SMC2, a factor of roughly 6.1. The variability in the CPU across the run in both methods reflects the stochastic occurrence of rejuvenation steps, which introduce random fluctuations in computing time.
As an additional benchmark, we evaluated eSMC2 against the Liu and West filter (Liu & West 2001), an alternative for joint sequential state-parameter estimation. Results in Appendix E demonstrate that eSMC2 achieves superior parameter accuracy and produces narrower, more stable credible intervals, particularly for correlated parameters.
Table 1: Comparison of goodness-of-fit and parameter estimates. Each method is evaluated based on CPU time in seconds (with standard deviation) and MAE (RMSE in parentheses) for daily incidence, effective reproduction number, transmission rate and model static parameters. Results are shown for both examples.
Method CPU (sd) MAE (RMSE)
Incidence ($\bm{\rho Z_{t}}$) $\bm{R_{\text{eff}}(t)}$ $\bm{\beta_{t}}$ $\bm{\widehat{\mathrm{E}}[\alpha|y_{1:T}]}$ $\bm{\widehat{\mathrm{E}}[\gamma|y_{1:T}]}$
@@
SMC2 3929 $\pm$ 433 4.652 (6.536) 0.266 (0.450) 0.032 (0.050) 0.072 (0.087) 0.049 (0.061)
eSMC2 639 $\pm$ 48 4.780 (6.697) 0.233 (0.386) 0.033 (0.053) 0.064 (0.078) 0.036 (0.046)

### v2

@@ 2.2 Ensemble Kalman Filter
$\widehat{p}_{\text{enkf}}^{N_{x}}(y_{t}|y_{1:t-1},\theta)=\widehat{\mathcal{N}}_{\text{unb}}\left(y_{t};~H\widehat{\mu}_{t|t-1},~H\widehat{\Sigma}_{t|t-1}H^{\top}+\widehat{V}_{t}\right).$ (17)
@@
Unlike in the case of the particle filter (Andrieu et al., 2010), the resulting posterior approximation $\widehat{p}_{\text{enkf}}^{N_{x}}(\theta|y_{1:t})\propto\widehat{p}_{\text{enkf}}^{N_{x}}(y_{1:t}|\theta)\,p(\theta)$, still does not target the true posterior (due to the Gaussian assumption). However, when the Gaussian assumption holds, the resulting posterior no longer depends on $N_{x}$, eliminating the finite-ensemble bias and yielding a stable target for Bayesian inference (Ghurye and Olkin, 1969; Drovandi et al., 2022). The EnKF procedure, including unbiased likelihood estimation, is outlined in Algorithm 1. This estimator can then be incorporated into a Metropolis–Hastings algorithm to target the sequence of posterior distributions $\{p(\theta),~p(\theta|y_{1}),~p(\theta|y_{1:2}),\ldots,~p(\theta|y_{1:T})\}$.
@@
Inputs: Observation $y_{1:T}$, Number of ensemble members $N_{x}$, Parameter vector $\theta$.
@@
4:  Compute the observation variance, ensemble forecast mean, and variance: $\widehat{V}_{t}$, $\widehat{\mu}_{t|t-1}$, and $\widehat{\Sigma}_{t|t-1}$.
@@
$\widehat{p}_{\text{enkf}}^{N_{x}}(y_{t}|y_{1:t-1},\theta)=\widehat{\mathcal{N}}_{\text{unb}}\left(y_{t};~H\widehat{\mu}_{t|t-1},~H\widehat{\Sigma}_{t|t-1}H^{\top}+\widehat{V}_{t}\right).$
@@ 2.3 Ensemble SMC2
The EnKF algorithm described previously assumes that the model parameters $\theta$ are known. In practice, these parameters are rarely known, and it is preferable to assign a distribution $p(\theta)$ that reflects prior knowledge. To account for this uncertainty, inference on $\theta$ can be performed by embedding the EnKF within an SMC sampler over the parameter space (Del Moral et al., 2006). This approach follows the general principle of SMC2 (Chopin et al., 2013), which traditionally employs a particle filter to estimate the likelihood. At each observation time $t$, the posterior distribution of the parameters given data $y_{1:t}$ satisfies the recursive relation
@@
As the incremental likelihood $p(y_{t}|y_{1:t-1},\theta)$ is intractable in a nonlinear system, at time $t$, the weight of the $\theta$-particle $\theta^{m}$ is updated using an estimator $\widehat{p}_{\text{enkf}}^{N_{x}}(y_{t}|y_{1:t-1},\theta^{m})$ obtained from the associated EnKF. This differs from related work of Wu et al. (2022), where the EnKF is used to construct proposal kernels in the SMC sampler but not as a likelihood estimator. As in standard SMC2, particle degeneracy may occur over time. To mitigate this, resampling (described in Appendix B) is triggered whenever the effective sample size (ESS) drops below $N_{\theta}/2$ (Chopin et al., 2013, 2020). This is followed by a jittering mechanism that rejuvenates the particles by proposing new candidates, which are then accepted or rejected through a few iterations of Particle Marginal Metropolis–Hastings (PMMH). Since the weighted particles already approximate $p(\theta|y_{1:t})$ (under the Gaussian assumption), applying a Metropolis-Hastings update to each one can only enhance this approximation (see Algorithm 2).
@@
$\vartheta(\theta^{*},\theta)=\min\left\{1,~\dfrac{\widehat{p}_{\text{enkf}}^{N_{x}}(y_{1:t}|\theta^{*})p(\theta^{*})}{\widehat{p}_{\text{enkf}}^{N_{x}}(y_{1:t}|\theta)p(\theta)}\times\dfrac{q(\theta|\theta^{*})}{q(\theta^{*}|\theta)}\right\}$
@@
The eSMC2 algorithm can be viewed as a variant of the standard SMC2, where the particle filter–based likelihood is replaced by an EnKF approximation in (18). From a practical perspective, this modification is expected to have only a slight effect on the posterior estimates of $\theta$ and $x_{1:t}$. The complete procedure is detailed in Algorithm 3, with a schematic overview in Fig 2.
To propagate parameter uncertainty in the hidden state estimation, we approximate the marginal posterior distribution of the latent states at time $t$, $p(x_{t}|y_{1:t})$, by integrating over a subset of parameter particles, $\{\theta^{j}\}_{j=1}^{N_{c}}\sim p(\theta|y_{1:T})$, drawn from Algorithm 3 output (Steyn et al., 2025). In this work, we set $N_{c}=100$, which in our numerical examples (Section 3) provides satisfactory coverage of the posterior. For each sampled $\theta^{j}$, an EnKF with $N_{x}$ ensemble members is run to approximate the filtering distribution $p(x_{t}|y_{1:t},\theta^{j})$, and the resulting trajectories are combined across all parameter samples to form an empirical approximation of the state posterior:
@@
It is important to note a key theoretical distinction between SMC2 and eSMC2. Standard SMC2 is “exact”: as the number of state and parameter particles $N_{x},N_{\theta}\to\infty$, it converges to the true posterior $p(\theta|y_{1:t})$. In contrast, eSMC2 is generally biased because the EnKF approximates the incremental likelihood $p(y_{t}|y_{1:t-1},\theta)$ via a Gaussian assumption. Nevertheless, the EnKF likelihood typically has low variance (Katzfuss et al., 2020), and unbiased corrections are possible under mild conditions (Drovandi et al., 2022). Empirical studies show that EnKF-based likelihoods yield accurate and stable inference for both linear and moderately nonlinear SSMs (Khalil et al., 2015; Drovandi et al., 2022). Thus, while not formally exact, eSMC2 provides practically accurate posterior estimates at much lower computational cost.
@@ 3.1 Inference on a simulated epidemic
To assess the performance of the proposed eSMC2 algorithm, we first conducted hree simulation studies using the SEIR model from Section 2. All algorithms were implemented in Python and executed on a desktop computer equipped with an Intel Core i7-1300H processor (3.40 GHz). Unless otherwise stated, simulations assumed a population size of N=500, [^0] N=500{,}000 with initial conditions $I_{0}=10$, $S_{0}=N-I_{0}$, and $E_{0}=R_{0}=0$. Daily incidence $y_{t}$ was generated according to the Poisson observation model in Eq (3), assuming perfect case reporting ($\rho=1$).
The three examples differ primarily in the temporal structure of the transmission rate $\beta_{t}$ and the number of observation points. Example 1 depicts a moderately time-varying epidemic with an initial surge followed by a mild resurgence, capturing short-term fluctuations in transmission. Example 2, by contrast, exhibits a smoother and more prolonged pattern that sustains infections for longer and decays more gradually. The latter setting was designed to emulate the dynamic observed in the 2022 U.S. mpox epidemic, where behavioral changes and intermittent control measures led to prolonged and uneven declines in case counts. Example 3 is motivated by epidemics such as COVID-19 or seasonal influenza that exhibit secondary infection waves, arising from imperfect immunity or the activation and deactivation of control policies. To capture this behavior, we employ an oscillatory transmission rate that generates recurrent epidemic waves of declining amplitude, providing a more challenging multi-wave setting in which to evaluate algorithmic performance.
@@
Synthetic data were generated over $T=60$ time points with time-varying transmission rate $\beta_{t}=0.3\cdot\exp\big(\sin(2\pi t/55)-t/80\big)$. The incubation and infectious periods were set to $\alpha=1/2$ and $\gamma=1/7$, respectively. Each state ensemble was initialized using truncated normal distributions: $S_{0}\sim\mathcal{TN}_{[0,\,\infty]}(5\times 10^{5},0.2^{2})$, $I_{0}\sim\mathcal{TN}_{[0,\,\infty]}(10,0.2^{2})$ and $E_{0}=R_{0}=0$ .The priors were specified as $\beta_{0}\sim\mathcal{N}(0.3,0.01^{2}),~\alpha\sim\mathcal{TN}_{[0,\,\infty]}(0.6,0.3^{2}),~\gamma\sim\mathcal{TN}_{[0,\,\infty]}(0.2,0.1^{2}),~\nu_{\beta}\sim\mathcal{U}(0,0.5).$
@@
Synthetic data were generated over $T=100$ time points with a time-varying transmission rate $\beta_{t}=0.5\cdot\exp\Big(-(t-15)^{2}/20^{2}\Big)+0.065$, and fixed parameters $\alpha=1/3$ and $\gamma=1/8$. The priors were specified as $\beta_{0}\sim\mathcal{N}(0.35,0.01^{2}),~\alpha\sim\mathcal{TN}_{[0,\,\infty]}(0.4,0.2^{2}),~\gamma\sim\mathcal{TN}_{[0,\,\infty]}(0.12,0.2^{2}),~\nu_{\beta}\sim\mathcal{U}(0,0.3).$ Initial state ensemble members were identical to Example 1.
Example 3.
Synthetic data were generated over $T=100$ time points with an oscillatory, multi-wave transmission rate $\beta_{t}=0.3\cdot\exp\big(\cos(2\pi t/70)-t/100\big)$, and fixed parameters $\alpha=1/2$ and $\gamma=1/6$. The priors were specified as $\beta_{0}\sim\mathcal{N}(0.8,0.01^{2}),~\alpha\sim\mathcal{TN}_{[0,\,\infty]}(0.6,0.3^{2}),~\gamma\sim\mathcal{TN}_{[0,\,\infty]}(0.17,0.05^{2}),~\nu_{\beta}\sim\mathcal{U}(0,1).$ A more informative prior was placed on $\gamma$ relative to Examples 1 and 2, as the oscillatory structure of $\beta_{t}$ introduces identifiability challenges that make the recovery rate harder to pin down from incidence data alone. Initial state ensemble members were identical to Examples 1 and 2.
For each simulation study, we compared the proposed eSMC2 algorithm with the standard SMC2 implementation that employs a bootstrap particle filter (BPF) in its inner layer (see Appendix A). In all cases, we used $N_{\theta}=1000$ parameter particles and $N_{x}=200$ ensemble members (or particles). For simplicity and comparability, the value of $N_{x}$ was kept fixed throughout the timeline. Parameter rejuvenation was triggered when the effective sample size of the parameter particles dropped below 50% of $N_{\theta}$, followed by five iterations of the PMMH kernel.
Filtered estimates in Figs 3,  4, and 5 correspond to a single representative run of each method. The results illustrate that both eSMC2 (blue) and SMC2 (orange) accurately capture the epidemic dynamics in Examples 1, 2, and 3. For both methods, the predicted incidence closely follows the observed data, with 95% credible intervals consistently encompassing the actual case counts, indicating reliable uncertainty quantification. Estimates of the time-varying transmission rate $\beta_{t}$ and the effective reproduction number $R_{\text{eff}}(t)$ are broadly consistent across methods and closely align with the ground truth, demonstrating that both approaches yield comparable assessments of the evolving transmission dynamics. The wider credible intervals for $R_{\text{eff}}(t)$ relative to the filtered incidence reflect two complementary sources of uncertainty: the roughness of the random walk prior on $\beta_{t}$, and parameter non-identifiability, particularly during the early epidemic phase when different combinations of $\beta_{t}$, $\alpha$, and $\gamma$ can yield indistinguishable incidence curves. This wider uncertainty is therefore an accurate signal that incidence data alone cannot sharply identify instantaneous transmission intensity. Smoother priors on $\beta_{t}$, such as integrated Brownian motion or higher-order random walks, could potentially mitigate this variability but were not considered here and remain an avenue for future work. The filtering estimates of the unobserved latent state also closely track the true latent trajectories, with their credible intervals providing appropriate coverage (see Appendix C). The filtering estimates of the unobserved latent state also closely track the true latent trajectories, with their credible intervals providing appropriate coverage (see Appendix C).
@@
Figs 6 and 7 and 8 illustrate the temporal evolution of the posterior means of the inferred parameters, along with their marginal posterior densities at the final time step, each summarizing five independent runs. Across runs, both SMC2 and eSMC2 accurately recover the ground truth. For the latency and recovery rates ($\alpha$ and $\gamma$), the posterior means obtained from both algorithms are closely aligned, with credible intervals narrowing steadily as more data are assimilated, indicating progressive information gain. The only notable exception is the diffusion parameter $\nu_{\beta}$ in Examples 2 and 3, where a slightly larger discrepancy is observed. This difference likely arises from the stochastic nature of the log-transmission process combined with the approximation introduced by the EnKF-based likelihood in eSMC2. Importantly, these differences do not introduce systematic bias or distort the inferred $\beta_{t}$ trajectory. The joint posterior contour plots (see Appendix C) further reveal strong agreement in posterior geometry and dependence structure: the contours for SMC2 and eSMC2 largely overlap, are centered near the true values, and display similar correlation patterns among parameters. The only visible difference appears along the $\nu_{\beta}$ dimension. Additional sensitivity analyses (Appendix D) assessing the effect of the number of state particles ($N_{x}$) and the reporting fraction ($\rho$) similarly demonstrate that posterior inferences are robust to ensemble size and exhibit only moderate variation when $\rho$ is correctly specified. The analyses also investigate the impact of non-informative (flat) priors, demonstrating that posterior estimates remain stable under weakly informative prior specifications.
@@
Results in Table 1 summarize the accuracy and computational cost of each method, averaged over five independent runs. The mean absolute error (MAE) and root mean squared error (RMSE) quantify discrepancies between the filtering means and the ground truth for both selected latent state ($Z_{t},~\beta_{t}$) and static parameters ($\alpha,~\gamma$). Across both examples, eSMC2 achieves error magnitudes closely matching those of SMC2, indicating that the ensemble approximation preserves inferential accuracy. In terms of computational effort, eSMC2 offers a substantial advantage. In Example 1, the average CPU time is approximately 467 seconds for eSMC2, compared with 2760 seconds for SMC2, corresponding to a reduction by a factor of about 4.87. Similarly, in Example 2, eSMC2 reduces average runtime to 639 seconds versus 3929 seconds for SMC2, a factor of roughly 6.1. In Example 3, eSMC2 completes in approximately 543 seconds on average, compared with 3766 seconds for SMC2, corresponding to a speedup factor of roughly 6.9. The variability in the CPU across the run in both methods reflects the stochastic occurrence of rejuvenation steps, which introduce random fluctuations in computing time.
Table 1: Comparison of goodness-of-fit and parameter estimates. Each method is evaluated based on CPU time in seconds (with standard deviation) and MAE (RMSE in parentheses) for daily incidence, effective reproduction number, transmission rate and model static parameters. Results are shown for all three examples.
Method CPU (sd) Incidence ($\bm{\rho Z_{t}}$) $\bm{R_{\text{eff}}(t)}$ $\bm{\beta_{t}}$ $\bm{\widehat{\mathrm{E}}[\alpha|y_{1:T}]}$ $\bm{\widehat{\mathrm{E}}[\gamma|y_{1:T}]}$
@@
[…]
Example 3
[…]
As an additional benchmark, we evaluated eSMC2 against the Liu and West filter (Liu and West, 2001) (using $10^{4}$ particles, runtime $\approx 9$ minutes) and a long PMCMC run ($10^{4}$ iterations, runtime $\approx 3$ hours) as a gold-standard reference, both serving as alternative approaches for joint sequential state-parameter estimation. Results provided in Appendix E demonstrate that eSMC2 achieves superior parameter accuracy and produces narrower, more stable credible intervals compared to the Liu and West filter, while remaining in close agreement with the PMCMC posterior distributions evaluated at the last time point.